OpenStudy (anonymous):

write an equation of a line in slope intercept form that is perpendicular to y=-4x-2 and passes through point (-16,-11)

OpenStudy (zale101):

Any ideas of where or how to start?

OpenStudy (anonymous):

You want to find the equation for a line that passes through the two points: (-16,-11) and (,). First of all, remember what the equation of a line is: y = mx+b Where: m is the slope, and b is the y-intercept First, let's find what m is, the slope of the line... The slope of a line is a measure of how fast the line "goes up" or "goes down". A large slope means the line goes up or down really fast (a very steep line). Small slopes means the line isn't very steep. A slope of zero means the line has no steepness at all; it is perfectly horizontal. For lines like these, the slope is always defined as "the change in y over the change in x" or, in equation form: So what we need now are the two points you gave that the line passes through. Let's call the first point you gave, (-16,-11), point #1, so the x and y numbers given will be called x1 and y1. Or, x1=-16 and y1=-11. Also, let's call the second point you gave, (,), point #2, so the x and y numbers here will be called x2 and y2. Or, x2= and y2=. Now, just plug the numbers into the formula for m above, like this: m= - -11 - -16 or... m= 11 16 or... m=11/16 So, we have the first piece to finding the equation of this line, and we can fill it into y=mx+b like this: y=11/16x+b Now, what about b, the y-intercept? To find b, think about what your (x,y) points mean: (-16,-11). When x of the line is -16, y of the line must be -11. (,). When x of the line is , y of the line must be . Because you said the line passes through each one of these two points, right? Now, look at our line's equation so far: y=11/16x+b. b is what we want, the 11/16 is already set and x and y are just two "free variables" sitting there. We can plug anything we want in for x and y here, but we want the equation for the line that specfically passes through the two points (-16,-11) and (,). So, why not plug in for x and y from one of our (x,y) points that we know the line passes through? This will allow us to solve for b for the particular line that passes through the two points you gave!. You can use either (x,y) point you want..the answer will be the same: (-16,-11). y=mx+b or -11=11/16 × -16+b, or solving for b: b=-11-(11/16)(-16). b=0. (,). y=mx+b or =11/16 × +b, or solving for b: b=-(11/16)(). b=0.